103rd BLM Thresholds Meeting (changes during YETS and system readiness)

Europe/Zurich
865/1-B03 (CERN)

865/1-B03

CERN

12
Show room on map
Anton Lechner (CERN), Belen Maria Salvachua Ferrando (CERN), Daniele Mirarchi (CERN), Sara Morales Vigo
Zoom Meeting ID
68180296658
Host
Sara Morales Vigo
Alternative host
Anton Lechner
Useful links
Join via phone
Zoom URL
  • B.Salvachua - Changes on the LHC BLM system during YETS
    • New BLM installations during EYETS 24-25
      • TWOCRYST, two new BLMs, one downstream splitting crystal (TCCS) and another one downstream precesion crystal (TCCP)
        • Not interlocked
        • Channel for TCCP crystal reused from BLM at TCSM empty slot
        • Installation and motivation for experiment in ECR...
      • BGC, installation of two BLMs to investigate possibility to use it for beam halo monitoring
        • Not interlocked
        • Installation and motivation in ECR...
      • RF cavities, relocation of three BLMs and installation of two additional ones around the cryo-module 1 - Beam 2
        • Were not interlocked and will not be now 
        • ECR under preparation
        • Cables now routed on the protection to shield from noise
      • TDIS 4L2, one BLM that had to be displaced by 66 cm
        • Modification of expert name to reflect the modification of dcum
        • No modification of thresholds
      • MKIA MKIB in IP8, removal and reinstallation of 4 BLMs in the same locations
        • No changes
      • LESS solenoid installation, removal and reinstallation of 2 BLMs in the same locations
        • No changes
      • Question from Cedric about timeline for LIC detectors
        • Belen says that Sevgi is performing LIC and IC simulations to verify conversion factors and signals 
        • Action: Check if some of the LIC detectors are already interlocked in injection areas, and what is the thresholds
      • Question from Daniele about dBLMs in IP3
        • Belen says that they have been installed, details about high and low gain channels will come
        • Logging already in place
        • The card was taken from SPS and cables from collimators
      • Question from Daniele about interlock on TWOCRYST BLMs
        • Concerned that if they are not interlocked, if there are problems with the reversal after MDs, it will not be noticed easily
        • Possibility to interlock BLMs and put thresholds to the maximum for situations like these
          • Decision -> This will be the case for the TWOCRYST BLMs

 

  • S.Morales - LHC BLM system readiness for 2025 first beam
    • Redundant 100Hz logging -> logged now by BLM per crate, also from Collimation
      • Build crontask to verify logging
    • Beam expected in 14 April
    • Latency tests: Check the note LHC-BLM-NOT-0008 with Daniele.
      • Latency in the ARC not tested
    • Cedric: Checklist in MPP
    • Cedric: in Friday there will be an update on the checklist

 

Next meeting 8th April - in the control room.

There are minutes attached to this event. Show them.